OpenStudy (anonymous):

f(x)=13/x-6 what is the domain of f

OpenStudy (mathstudent55):

Oh, i see. you started a new post for this question.

OpenStudy (usukidoll):

the domain is the x-axis... but this time we have a restriction...

OpenStudy (mathstudent55):

Do you mean \(f(x) = \dfrac{13}{x - 6} \) ?

OpenStudy (usukidoll):

I'm sure that's right... because there is a restriction for that function..

OpenStudy (anonymous):

yes

OpenStudy (usukidoll):

for what value of x makes the entire function undefined ?

OpenStudy (mathstudent55):

The domain is the set of values that x can have. A fraction can never have a denominator of zero. If the denominator is x - 6, then x - 6 cannot be zero.

OpenStudy (usukidoll):

it's similar to solving x - 6 = 0

OpenStudy (mathstudent55):

To find what value makes the denominator zero, solve the simple equation: x - 6 = 0 What is x?

OpenStudy (usukidoll):

what do we need to do to have x by itself?

OpenStudy (anonymous):

0

OpenStudy (usukidoll):

no.. we are solving for x for x -6=0

OpenStudy (mathstudent55):

No. If x = 0, x - 6 = 0 - 6 = -6. That makes a denominator of -6 which is fine.

OpenStudy (mathstudent55):

How do you solve the equation x - 6 = 0 for x?

OpenStudy (usukidoll):

how do we get x by itself? there must be some arithmetic somewhere x-6=0

OpenStudy (mathstudent55):

Since you want to isolate x, and 6 is being subtracted from x, you add 6 to both sides.

OpenStudy (mathstudent55):